# Unbiased Sliced Wasserstein Kernels for High-Quality Audio Captioning

### v1

@@ 3.2 Audio captioning with the Unbiased SW-RBF kernel framework
Positional encoding for USW-RBF kernel. Given a pair of audio and ground-truth caption is denoted as $(x,y)$, the hidden representation of audio, extracted from the penultimate layer of the audio encoder, is denoted as $Z_{x}=[z_{x}^{1},...,z_{x}^{N}]$, where $z_{x}^{i}\in\mathbb{R}^{d}$, and the hidden representation of ground-truth caption conditioning on the audio, extracted from the penultimate layer of the decoder, is denoted as $Z_{y}=[z_{y}^{1},...,z_{y}^{M}]$ where $z_{y}^{j}\in\mathbb{R}^{d}$. Although the USW-RBF is effective in measuring the similarity between two sets of vectors, the order of vectors within a set is not taken into account when computing the sliced Wasserstein distance. More importantly, the order of vectors within a set contains the temporal information between them, which is crucial for audio and language modality. To preserve the temporal information, we define the temporal-information preserving vector as follows
$\phi_{x}^{n}=concat(z_{x}^{n},pos(n))$ (11)
where $n$-th denotes the position of vector $z^{n}_{x}\in\mathbb{R}^{d}$ in a sequence of vector $Z_{x}\in\mathbb{R}^{N\times d}$, and $pos(n)\in\mathbb{R}^{k}$ is the corresponding positional embedding vector. there are two popular positional embedding functions: absolute positional embedding (Vaswani et al., 2017) and rotary positional embedding functions (Su et al., 2024). We redefine $Z_{x}=[\phi_{x}^{1},\ldots,\phi_{x}^{N}]$ and $Z_{y}=[\phi_{y}^{1},\ldots,\phi_{y}^{M}]$ respectively.
Training with the USW-RBF kernel. We assume that $N=M$, two projected-one dimensional sequences $a_{\psi}=[a_{1},...,a_{N}]$ and $b_{\psi}=[b_{1},...,b_{N}]$, where $a_{i}=\psi^{\top}\phi^{i}_{x}$ and $b_{j}=\psi^{\top}\phi^{j}_{y}$. We denote the $\sigma_{1}:[[N]]\to[[N]]$ and $\sigma_{2}:[[N]]\to[[N]]$ as two sorted permutation mapping of $a_{\psi}$ and $b_{\psi}$ in turn. Let denote the projection vector $\psi=concat(\psi_{1},\psi_{2})$ is the concatenation of two vectors $\psi_{1}\in\mathbb{R}^{d}$ and $\psi_{2}\in\mathbb{R}^{k}$. Now, we define the temporal-similarity score based USW-RBF with $p=2$:
$\mathcal{UK}_{\gamma}(Z_{x},Z_{y};2)$ (12)
$=\mathbb{E}_{\psi\sim\mathcal{U}(\mathbb{S}^{d+k-1})}\left[\exp\left(-\gamma\sum_{i=1}^{N}(a_{\sigma_{\psi,1}(i)}-b_{\sigma_{\psi,2}(i)})^{2}\right)\right]$
$=\mathbb{E}_{\psi\sim\mathcal{U}(\mathbb{S}^{d+k-1})}\left[\exp\left(-\gamma\sum_{i}^{N}\left[\left(\underbrace{\psi_{1}^{\top}z_{x}^{\sigma_{1}(i)}-\psi_{1}^{\top}z_{y}^{\sigma_{2}(i)}}_{\text{\clap{$K_{\psi,1}$~{}}}}\right.\right.\right.\right.$
$\left.\left.\left.\left.+\underbrace{\psi_{2}^{\top}pos(\sigma_{1}(i))-\psi_{2}^{\top}pos(\sigma_{2}(i))}_{\text{\clap{$K_{\psi,2}$~{}}}}\right)^{2}\right]\right)\right]$
@@
$\mathcal{L}=\mathcal{L}_{MLE}(x,y)+\mathcal{UK}_{\gamma}(Z_{x},Z_{y};2).$ (13)
Inference stage. As extensively discussed in the literature, likelihood decoding is suffering from exposure bias (An et al., 2022; Su et al., 2022). A solution is to utilize stochastic decoding, such as top-k or nucleus sampling  (Holtzman et al., 2020)methods, to mitigate the harmful effect of exposure bias (Arora et al., 2022b). We propose to leverage the temporal-similarity score based on the USW-RBF between the latent representation of audio and generated captions as a decoding criterion. As demonstrated in the Figure 1, the pretrained audio captioning model generates $\mathcal{B}$ candidate captions by stochastic decoding methods, and the most likely caption is chosen as follows
$\mathbf{y^{*}}=\operatorname*{arg\,max}_{\mathbf{y_{i}}\in\mathcal{B}}\{(1-\alpha)p(\mathbf{y_{i}}|x)+\alpha.\mathcal{UK}_{\gamma}(Z_{x},Z_{y};2)$ (14)
where $Z_{x},Z_{\mathbf{y}_{i}}$ denote the latent representation of audio and generated captions outputted from the encoder and decoder models, respectively. The coefficient $10<\alpha<$ 1 is set to $0.5$ in the most case. The first term of the decoding objective is the likelihood score of a generated caption, which measures the confidence of the audio captioning model. The second term measures the similarity in terms of the latent representation of audio and generated captions.
@@ 4 Related Work
Audio captioning. The audio captioning task can be formulated as a conditional text generation task, therefore, the prior works utilize the maximum likelihood estimation method to train audio captioning models (Mei et al., 2021, 2024; Sun et al., 2023; Kim et al., 2022; Deshmukh et al., 2023). There are two popular architectures for audio captioning models: encoder-decoder architecture (Mei et al., 2024; Kim et al., 2024) and prefix-tuning architecture (Deshmukh et al., 2023; Kim et al., 2023). Although both architectures are effective in generating plausible captions, they suffer from the inherent weakness of the MLE training method: exposure bias. Some recent works deal with exposure bias by leveraging a regularization (Zhang et al., 2023; Deshmukh et al., 2024), contrastive loss. The contrastive regularization can slightly remedy the exposure bias issue for audio captioning models. Another technique to combat with exposure bias is to utilize stochastic decoding methods (Arora et al., 2022a).  (Su et al., 2022) proposed a contrastive search framework with stochastic decoding methods to alleviate text degeneration for conditional text generation. The contrastive search framework is yet successful to deal with exposure bias for text generation, it can not be directly applied for audio captioning task. The reason is that the contrastive score is not able to take temporal information of acoustic and linguistic features into account. To deal with the shortcomings of the contrastive framework, we develop a new framework, called ACUS, which can handle the temporal information between acoustics and linguistic modalities when measuring the similarity score and alleviate exposure bias at the inference stage for audio captioning.
Wasserstein distance. Wasserstein distance is a metric to measure the discrepancy between two distributions. There are enormous applications of the Wasserstein distance for multimodal learning, such as audio-text retrieval (Luong et al., 2024), multimodal representation learning (Tsai et al., 2019), and multimodal alginment (Lee et al., 2019). The prior work (Su & Hua, 2017) proposed an order-preserving Wasserstein distance between sequences by incorporating a soft-monotonic alignment prior for optimal matching, however, it still suffers from dimensionality curse and a strict monotonic alignment across modalities. Although the Wasserstein distance is capable of measuring the cross-modality distance, it suffers from the dimensionality curse. In this work, we develop the USW-RBF kernel equipped with positional encoding to deal with the dimensionality curse and the strict monotonic alignment issue of measuring cross-modal similarity for audio captioning.
@@ 5 Experiments
Table 1: The quantitative evaluation of proposed method with baselines using objective metrics on AudioCaps and Clotho datasets. The ACUS and contrastive frameworks utilize stochastic decoding methods during the inference stage, therefore, we report the average performance and standard deviation for these methods.
Dataset Method METEOR ROUGE_L CIDEr SPICE SPIDEr
AudioCaps ACT 0.222 0.468 0.679 0.160 0.420
LHDFF 0.232 0.483 0.680 0.171 0.426
@@
Our method $\textbf{0.262}\pm\textbf{0.001}$ $\textbf{0.509}\pm\textbf{0.001}$ $\textbf{0.807}\pm\textbf{0.003}$ $\textbf{0.192}\pm\textbf{0.001}$ $\textbf{0.5}\pm\textbf{0.002}$
@@
Our method $\textbf{0.186}\pm\textbf{0.001}$ $\textbf{0.38}\pm\textbf{0.001}$ $\textbf{0.419}\pm\textbf{0.004}$ $\textbf{0.133}\pm\textbf{0.001}$ $\textbf{0.275}\pm\textbf{0.003}$
Model Decoding METEOR ROUGE_L CIDEr SPICE SPIDEr
ACT Beam(k=5) 0.222 0.468 0.679 0.160 0.420
Enclap Beam(k=5) 0.254 0.5 0.77 0.186 0.48
We design experiments to demonstrate the effectiveness of our proposed method in mitigating exposure bias in the audio captioning task. We conduct quantitative experiments on two datasets: Audiocaps (Kim et al., 2019) and Clotho (Drossos et al., 2020) to answer the question of whether our proposed method is capable of alleviating exposure bias in the audio captioning task. We further conduct qualitative experiments on audio-text retrieval tasks and subjective evaluation to show the high-quality of generated captions. Finally, we perform ablation studies on the choice of similarity metric and positional embedding techniques. The ablation studies show that the proposed metric outperforms both Wasserstein distance,  DTW, and soft-DTW in measuring the similarity between latent representation of audio and generated captions. These studies also show that rotary positional embedding is the most well-suited positional embedding technique for incorporating temporal information for audio-captioning.
Evaluation metrics. We evaluate baselines and two backbone models, Enclap and ACT, for our proposed framework by widely used evaluation metrics for audio captioning, including METEOR (Banerjee & Lavie, 2005), ROUGE-L (Lin, 2004), CIDEr (Vedantam et al., 2014), SPICE (Anderson et al., 2016), and SPIDEr (Liu et al., 2016). In addition, we evaluate the quality of generated audio captions by performing a text-to-audio retrieval task leveraging the pretrained CLAP (Wu et al., 2023) model. If a generated caption and a given audio are highly similar to each other, the CLAP model is able to retrieve the audio by using the generated caption. We further measure the lexical diversity and caption length in generated captions to measure the degeneration of captions. We also conduct a subjective evaluation to evaluate the quality of generated captions in terms of discretiveness, correctness, and fluency.
Baselines. We compare against all state-of-the-art audio captioning models on Audiocaps and Clotho datasets. The ACT (Mei et al., 2021) audio captioning model leverages a vision transformer encoder pretrained on the AudioSet (Gemmeke et al., 2017) dataset for sound-event classification. LHDFF (Sun et al., 2023) utilizes residual the PANNs encoder to fuse low and high dimensional features in Mel-spectrogram. CNN14-GPT2 (Kim et al., 2023), Pengi (Deshmukh et al., 2023), and RECAP (Ghosh et al., 2023) apply prefix-tuning method for the pretrained GPT2 (Radford et al., 2019). AL-MixGen (Kim et al., 2022) leverages the ACT backbone trained using audio-language mixup augmentation and test-time augmentation at the inference phase. Wavcaps (Mei et al., 2024) is the HTSAT-BART model (Chen et al., 2022b) fine-tuned on numerous weakly-labeled data which is generated by using large language models. We choose a subset of models evaluated on the Clotho dataset without complex training methods, such as ensemble training, to ensure a fair comparison. The CLIP-AAC (Chen et al., 2022a), MAAC (Ye et al., 2021), P-LocalAFT(Xiao et al., 2022), and Graph-AC (Xiao et al., 2023) are the baselines evaluated on Clotho dataset.
5.1 Implementation details
@@
ACT backbone. We follow the original settings in (Mei et al., 2021) to train the audio captioning transformer (ACT) backbone on the AudioCaps dataset. We use a batch size of 32 and warm up for five epochs before reaching the peak learning rate at $lr=1e^{-4}$. We use the training objective function in Equation (13) with training coefficient $\alpha=0.1$ and the bandwidth for the temporal-similarity metric $\gamma=2.5$. We also perform a grid search for stochastic decoding methods at the inference state to choose the best hyperparameters for each stochastic decoding method, $p=\{0.5,0.6,0.7,0.8,0.9\}$ for top-p sampling, $k=\{3,4,5\}$ for top-k sampling, and $temp=\{1.1,1.2,1.3,1.4,1.5\}$ for temperature sampling. The best results with optimal decoding hyperparameters are reported in Table 2.
5.2 Quantitative Experiments
To assess the performance of our proposed method for audio captioning, we performed quantitative experiments on Audiocaps and Clotho. The experimental results are shown in the Table. 1. All baseline models utilize deterministic decoding methods, the beam search decoding, therefore their performance is not variant in each evaluation. On the other hand, the contrastive method and our framework utilize stochastic decoding methods, such as the nucleus and top-k samplings, thus its performance varies for each evaluation. To make a fair comparison, we evaluate both our framework and contrastive method 5 times and report the average performance and standard deviation. It is clear to see that our proposed method outperforms all baseline models in terms of automated metrics on the AudioCaps test set. Specifically, our proposed framework significantly improves the quality of generated captions for the Enclap backbone model. There is a significant improvement regarding the statistical metrics SPICE, METEOR, CIDEr, and ROUGE-L. These results prove that our proposed method is able to mitigate the exposure bias for audio captioning models during inference. Furthermore, there is a significant performance gain regarding the SPICE score, from $0.186$ to $0.192$. Since the SPICE score captures the semantic similarity between generated and ground-truth captions, the proposed method is able to generate better semantically similar captions with reference. A similar improvement regarding objective metrics is observed for the Clotho dataset. The improvement is insignificant due to the diversity of reference captions in the Clotho dataset for automated metrics like $\text{ROUGE}_{L}$ and CIDEr that rely on measuring statistical overlap between predicted and reference captions.
In Table 2, we conducted the experiment on the diverse audio captioning backbones, the Enclap and ACT models, for the proposed method. The Enclap model is a encoder-decoder model which consists of a pretrained audio encoder from the CLAP model (Wu et al., 2023) and a pretrained BART decoder model. The ACT model is also a encoder-decoder model, which includes a vision transformer encoder pretrained on the AudioSet dataset and a transformer decoder model. The performance of backbone models with beam search decoding is substantially enhanced by our proposed approach when decoded with stochastic decoding techniques. The nucleus sampling technique with our method achieves the highest performance gain for both backbone models, while the stochastic decoding with temperature shows a little improvement. Especially, there is a slight drop in the CIDEr metric using stochastic decoding with temperature. The experimental results show the importance of controlling stochasticness when decoding to mitigate exposure bias. We also carry out ablation studies for choosing hyperparameters for stochastic decoding methods using our framework, and the results are reported in the Appendix A.3.
5.3 Qualitative Experiments

### v2

@@ 3.2 Audio captioning with the Unbiased SW-RBF kernel framework
Positional encoding for USW-RBF kernel. Given a pair of audio and ground-truth caption is denoted as $(\mathbf{x},\mathbf{y})$, the hidden representation of audio, extracted from the penultimate layer of the audio encoder, is denoted as $Z_{\mathbf{x}}=[z_{\mathbf{x}}^{1},...,z_{\mathbf{x}}^{N}]$, where $z_{\mathbf{x}}^{i}\in\mathbb{R}^{d}$, and the hidden representation of ground-truth caption conditioning on the audio, extracted from the penultimate layer of the decoder, is denoted as $Z_{\mathbf{y}}=[z_{\mathbf{y}}^{1},...,z_{\mathbf{y}}^{M}]$ where $z_{\mathbf{y}}^{j}\in\mathbb{R}^{d}$. Although the USW-RBF is effective in measuring the similarity between two sets of vectors, the order of vectors within a set is not taken into account when computing the sliced Wasserstein distance. More importantly, the order of vectors within a set contains the temporal information between them, which is crucial for audio and language modality. To preserve the temporal information, we define the temporal-information preserving vector as follows
$\phi_{x}^{n}=\operatorname{concat}(z_{\mathbf{x}}^{n},\operatorname{pos}(n))$ (11)
where $n$ denotes the position of vector $z^{n}_{\mathbf{x}}\in\mathbb{R}^{d}$ in a sequence of vector $Z_{\mathbf{x}}\in\mathbb{R}^{N\times d}$, and $\operatorname{pos}(n)\in\mathbb{R}^{k}$ is the corresponding positional embedding vector. there are two popular positional embedding functions: absolute positional embedding Vaswani et al. (2017) and rotary positional embedding functions (Su et al., 2024). We redefine $Z_{\mathbf{x}}=[\phi_{\mathbf{x}}^{1},\ldots,\phi_{\mathbf{x}}^{N}]$ and $Z_{\mathbf{y}}=[\phi_{\mathbf{y}}^{1},\ldots,\phi_{\mathbf{y}}^{M}]$ respectively.
Training with the USW-RBF kernel. We assume that $N=M$, two projected-one dimensional sequences $a_{\psi}=[a_{1},...,a_{N}]$ and $b_{\psi}=[b_{1},...,b_{N}]$, where $a_{i}=\psi^{\top}\phi^{i}_{\mathbf{x}}$ and $b_{j}=\psi^{\top}\phi^{j}_{\mathbf{y}}$. We denote the $\sigma_{1}:[[N]]\to[[N]]$ and $\sigma_{2}:[[N]]\to[[N]]$ as two sorted permutation mappings of $a_{\psi}$ and $b_{\psi}$ in turn. Let $\psi=\operatorname{concat}(\psi_{1},\psi_{2})$ denote the projection vector which is the concatenation of two vectors $\psi_{1}\in\mathbb{R}^{d}$ and $\psi_{2}\in\mathbb{R}^{k}$. Now, we define the temporal-similarity score based USW-RBF with $p=2$:
$\mathcal{UK}_{\gamma}(Z_{\mathbf{x}},Z_{\mathbf{y}};2)=\mathbb{E}_{\psi\sim\mathcal{U}(\mathbb{S}^{d+k-1})}\left[\exp\left(-\gamma\sum_{i=1}^{N}(a_{\sigma_{\psi,l}(i)}-b_{\sigma_{\psi,l}(i)})^{2}\right)\right]$ (12)
$=$ $\mathbb{E}_{\psi\sim\mathcal{U}(\mathbb{S}^{d+k-1})}\left[\exp\left(-\gamma\sum_{i}^{N}\left[\left(\underbrace{\psi_{1}^{\top}z_{\mathbf{x}}^{\sigma_{1}(i)}-\psi_{1}^{\top}z_{\mathbf{y}}^{\sigma_{2}(i)}}_{\text{\clap{$K_{\psi,1}$~}}}+\underbrace{\psi_{2}^{\top}pos(\sigma_{1}(i))-\psi_{2}^{\top}pos(\sigma_{2}(i))}_{\text{\clap{$K_{\psi,2}$~}}}\right)^{2}\right]\right)\right]$
@@
$\mathcal{L}=\mathcal{L}_{MLE}(\mathbf{x},\mathbf{y})+\mathcal{UK}_{\gamma}(Z_{\mathbf{x}},Z_{\mathbf{y}};2).$ (13)
Inference stage. As extensively discussed in the literature, likelihood decoding is suffering from exposure bias (An et al., 2022; Su et al., 2022). A solution is to utilize stochastic decoding, such as top-k or nucleus sampling  (Holtzman et al., 2020) methods, to mitigate the harmful effect of exposure bias (Arora et al., 2022b). We propose to leverage the temporal-similarity score based on the USW-RBF between the latent representation of audio and generated captions as a decoding criterion. As demonstrated in Figure 1, the pretrained audio captioning model generates $\mathcal{B}$ candidate captions by stochastic decoding methods, and the most likely caption is chosen as follows
$\mathbf{y^{*}}=\operatorname*{arg\,max}_{\mathbf{y}\in\mathcal{B}}\{p(\mathbf{y}|x)+\mathcal{UK}_{\gamma}(Z_{\mathbf{x}},Z_{\mathbf{y}};2)\}$ (14)
where $Z_{\mathbf{x}},Z_{\mathbf{y}_{i}}$ denote the latent representation of audio and generated captions outputted from the encoder and decoder models, respectively. The first term of the decoding objective is the likelihood score of a generated caption, which measures the confidence of the audio captioning model. The second term measures the similarity in terms of the latent representation of audio and generated captions.
@@ 4 Related Work
Audio captioning. The audio captioning task can be formulated as a conditional text generation task, therefore, the prior works utilize the maximum likelihood estimation method to train audio captioning models (Mei et al., 2021, 2024; Sun et al., 2023; Kim et al., 2022; Deshmukh et al., 2023). There are two popular architectures for audio captioning models: encoder-decoder architecture Mei et al. (2024); Kim et al. (2024) and prefix-tuning architecture (Deshmukh et al., 2023; Kim et al., 2023). Although both architectures are effective in generating plausible captions, they suffer from the inherent weakness of the MLE training method: exposure bias. Some recent works deal with exposure bias by leveraging a regularization (Zhang et al., 2023; Deshmukh et al., 2024), such as contrastive loss. The contrastive regularization can slightly remedy the exposure bias issue for audio captioning models. Another technique to deal with exposure bias is to utilize stochastic decoding methods (Arora et al., 2022a).  (Su et al., 2022) proposed a contrastive search framework with stochastic decoding methods to alleviate text degeneration for conditional text generation. Although the contrastive search framework is successful to deal with exposure bias for text generation, it can not be directly applied for audio captioning task. The reason is that the contrastive score is not able to take temporal information of acoustic and linguistic features into account. To deal with the shortcomings of the contrastive framework, we develop a new framework, called ACUS, which can handle the temporal information between acoustics and linguistic modalities when measuring the similarity score and alleviate exposure bias at the inference stage for audio captioning.
Wasserstein distance. Wasserstein distance is a metric to measure the discrepancy between two distributions. There are many applications of the Wasserstein distance for multimodal learning, such as audio-text retrieval (Luong et al., 2024), multimodal representation learning (Tsai et al., 2019), and multimodal alignment (Lee et al., 2019). The prior work (Su and Hua, 2017) proposed an order-preserving Wasserstein distance between sequences by incorporating a soft-monotonic alignment prior for optimal matching, however, it still suffers from dimensionality curse and a strict monotonic alignment across modalities. Although the Wasserstein distance is capable of measuring the cross-modality distance, it suffers from the dimensionality curse. In this work, we develop the USW-RBF kernel equipped with positional encoding to deal with the dimensionality curse and the strict monotonic alignment issue of measuring cross-modal similarity for audio captioning.
@@ 5 Experiments
We design experiments to demonstrate the effectiveness of our proposed method in mitigating exposure bias in the audio captioning task. We conduct quantitative experiments on two datasets: Audiocaps (Kim et al., 2019) and Clotho (Drossos et al., 2020) to answer the question of whether our proposed method is capable of alleviating exposure bias in the audio captioning task. We further conduct qualitative experiments on audio-text retrieval tasks and subjective evaluation to show the high-quality of generated captions. We further conduct experiments on two audio reasoning benchmarks, the CompA-R test (Ghosh et al., 2024) and the MMAU test mini benchmarks (Sakshi et al., 2025), to demonstrate the generalizability of our USW-RBF kernel to a broad range of cross-modal audio-text tasks. The ablation studies regarding the choice of similarity metrics, positional embedding techniques, efficiency and effectiveness trade-off, and hyper-parameter tuning for the USW-RBF kernel can be found in Appendix. 8. Baselines and implementation details can be found in Appendix A.2. The code of our ACUS framework is released in https://github.com/v-manhlt3/ACUS
Evaluation metrics. We evaluate baselines and two backbone models, Enclap and ACT, for our proposed framework by widely used evaluation metrics for audio captioning, including METEOR (Banerjee and Lavie, 2005), ROUGE-L (Lin, 2004), CIDEr (Vedantam et al., 2014), SPICE (Anderson et al., 2016), and SPIDEr (Liu et al., 2016). In addition, we evaluate the quality of generated audio captions by performing a text-to-audio retrieval task leveraging the pretrained CLAP (Wu et al., 2023) model. If a generated caption and a given audio are highly similar to each other, the CLAP model is able to retrieve the audio by using the generated caption. We further measure the lexical diversity and caption length in generated captions to measure the degeneration of captions. We also conduct a subjective evaluation to evaluate the quality of generated captions in terms of descriptiveness, correctness, and fluency.
5.1 Quantitative Experiments
Table 1: The quantitative evaluation of the proposed method with baselines using objective metrics on AudioCaps and Clotho datasets. The ACUS and contrastive frameworks utilize stochastic decoding methods during the inference stage, therefore, we report the average performance and standard deviation for these methods. ** denotes the reproduced results from public source code.
Dataset Method METEOR ROUGE-L CIDEr SPICE SPIDEr
AudioCaps LHDFF 0.232 0.483 0.680 0.171 0.426
@@ 5.1 Quantitative Experiments
ACUS(ours) $\textbf{0.262}\pm\textbf{0.001}$ $0.509\pm 0.001$ $\textbf{0.807}\pm\textbf{0.003}$ $\textbf{0.192}\pm\textbf{0.001}$ $\textbf{0.5}\pm\textbf{0.002}$
@@
ACUS(ours) $\textbf{0.186}\pm\textbf{0.001}$ $\textbf{0.38}\pm\textbf{0.001}$ $\textbf{0.419}\pm\textbf{0.004}$ $\textbf{0.133}\pm\textbf{0.001}$ $\textbf{0.275}\pm\textbf{0.003}$
To assess the performance of our proposed method for audio captioning, we performed quantitative experiments on Audiocaps and Clotho. The experimental results are shown in the Table. 1. All baseline models utilize deterministic decoding methods, the beam search decoding, therefore their performance is not variant in each evaluation. On the other hand, the contrastive method and our framework utilize stochastic decoding methods, such as the nucleus and top-k samplings, thus their performance varies for each evaluation. To make a fair comparison, we evaluate both our framework and the contrastive method 5 times and report the average performance and standard deviation. It is clear to see that our proposed method outperforms all baseline models across the majority of automated evaluation metrics, with the exception of the ROUGE-L metric, on the AudioCaps test set. Specifically, our proposed framework significantly improves the quality of generated captions for the Enclap backbone model. There is a significant improvement regarding the statistical metrics SPICE, METEOR, and CIDEr. These results demonstrate that our proposed method is able to mitigate the exposure bias for audio captioning models during inference. Furthermore, there is a significant performance gain regarding the SPICE score, from $0.186$ to $0.192$. Since the SPICE score captures the semantic similarity between generated and ground-truth captions, the proposed method is able to generate better semantically similar captions with reference. A similar improvement regarding objective metrics is observed for the Clotho dataset. The improvement is insignificant due to the diversity of reference captions in the Clotho dataset for automated metrics like ROUGE-L and CIDEr that rely on measuring statistical overlap between predicted and reference captions.
5.2 Qualitative Experiments
@@
Human evaluation. We conduct a human evaluation to better assess the quality of generated captions. We randomly choose 50 audios from AudioCaps and Clotho test data. Captions are generated for each audio by using different methods: maximum likelihood estimation (MLE), contrastive framework, and the ACUS framework. The MLE method utilizes a deterministic decoding method, beam search with a beam size of 5, while contrastive learning and the proposed method utilize a stochastic decoding method, top-p sampling with $p=0.7$ to generate 30 candidate captions. The most suitable caption is chosen based on Equation (5) for contrastive learning and Equation (14) for the proposed method. We recruit five annotators, who are asked to independently assess the quality of a given caption following a 5-point Likert scale for three aspects: descriptiveness, correctness, and fluency.
Table 3 shows the human valuation results on three aspects for Audiocaps and Clotho datasets. The inter-annotator agreement is shown in the last row measured by the Fleiss Kappa score (Fleiss, 1971). On both datasets, our method is capable of generating more descriptive and correct captions compared to baseline models trained with MLE and contrastive learning objectives. Also, all generated captions are more fluent than human-written captions. The rationale behind it is that humans focus more on audio content rather than fluency. On the other hand, audio captioning models leverage pretrained language models as the decoder, therefore, they can generate coherence captions; however, they tend to focus less on accurately describing audio content. The qualitative examples can be found in Appendix A.5.
5.3 Generalizability to audio reasoning tasks
Table 4: The comparison of USW-RBF kernel with contrastive learning metric for audio reasoning tasks on two benchmarks: CompA-R-test and MMAU test mini.
Method CompA-R-test (GPT4-o-score) MMAU test mini (Accuracy)
Clarity Correctness Engagement Average Sound Music Speech Average
GAMA 4.3 3.9 3.9 4.0 36.04 34.53 19.52 30.1
GAMA w/ CL 4.4 4.0 3.9 4.1 37.53 32.93 21.02 30.49
GAMA w/ USW-RBF 4.5 4.2 4.1 4.3 43.54 33.23 25.53 34.10
Table 5: The temporal sound event reasoning on the MMAU test mini benchmark. TER and ESR are temporal event reasoning and event-based sound reasoning questions, respectively.
Method MMAU test mini
TER ESR
GAMA 16.67 29.17
GAMA w/ CL 20.83 31.25
GAMA w/ USW-RBF 31.25 39.58
We extend the USW-RBF kernel to audio reasoning tasks to examine the generalizability of our proposed kernel to handle acoustic and linguistic alignment. We utilize the GAMA (Ghosh et al., 2024) model, which published both its pretrained parameters and instruction fine-tuning data, as a baseline model and then finetune the base GAMA model using the objective function described in Eq. 13. We compare our USW-RBF kernel with the contrastive learning metric for enhancing audio reasoning abilities of the GAMA model on two benchmarks: CompA-R-test Ghosh et al. (2024) and the MMAU test mini benchmark (Sakshi et al., 2025). The experimental results are shown in the Table. 4. We use the GPT4-o score (Ghosh et al., 2024) to benchmark the performance of the USW-RBF, comparing with the contrastive learning metric on the CompaA-R-test benchmark. The GPT4-o score evaluates three dimensions of reasoning responses: clarity, correctness, and engagement. Furthermore, we benchmark the performance baseline methods and our USW-RBF kernel by the accuracy metric on the MMAU test mini benchmark, which consists of reasoning questions for sound, music, and speech. Our kernel metric outperforms both the MLE and contrastive learning methods in terms of enhancing the clarity, correctness, and engagement of the GAMA model’s responses. Our method also increases the average accuracy of the base model from $30.1\%$ to $34.10\%$ on the MMAU test mini benchmark. The results in Table. 5 also show that our kernel metric is capable of improving the temporal event reasoning ability of large audio language models.
5.4 Ablation study